OpenStudy (anonymous):

\[x^3-y^3=xy+61\] \[x,y \in \mathbb{N}\]

OpenStudy (anonymous):

what restrictions are there for \(x\) and \(y\) ?

OpenStudy (anonymous):

they both should be natural numbers :P

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

haha nice !!

OpenStudy (anonymous):

and they should obviously satisfy the eqn

OpenStudy (anonymous):

nice man :D

OpenStudy (anonymous):

i mean \(x-y\) must be a limited natural number ... why?

OpenStudy (anonymous):

whay x can not be very greater than y ?

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

if x>>y then LHS will be very greater than RHS due exponents... 3(LHS) in comparison with 2(RHS)

OpenStudy (anonymous):

*due to exponents

OpenStudy (anonymous):

so let\[x-y=a\]a must be a restricted number equation becomes to :\[y^2(3a-1)+y(3a^2-2)+a^3-61=0\]

OpenStudy (anonymous):

y is positive and if a>something then LHS of later equation >0 and there is no solution so whats that something now?

OpenStudy (anonymous):

x and y both can be any numbers as long as they satisfy the equation

OpenStudy (anonymous):

we want to find all possible values of x and y that satisfies the equation

OpenStudy (anonymous):

if a>4 there is no solution for the later equation am i right?

OpenStudy (anonymous):

sorry man i am not getting it :(

OpenStudy (anonymous):

np man

OpenStudy (anonymous):

even we can tell if a>3 there is no solution for the later equation

OpenStudy (anonymous):

so \[a=1,2,3\]just from \(a=1\) we get the solution \((x,y)=(6,5)\).

OpenStudy (anonymous):

and there is no solution from \(a=2,3\)

hartnn (hartnn):

how did u get y(3a^2 - 2) ?? i m getting y(3a^2-a).....

OpenStudy (anonymous):

sorry thats right\[y^2(3a-1)+y(3a^2-a)+a^3-61=0\]

hartnn (hartnn):

a>3 no solution.....does it have something to do with discriminant of that equation?how did u get a>3 ,no solution?

OpenStudy (anonymous):

if a>3 all the coefficients of quadratic will be positive ... y is a positive number too...so the LHS became a positive expression

OpenStudy (anonymous):

and of course we workin with natural numbers here

hartnn (hartnn):

yes,i got that....thanks :)

OpenStudy (anonymous):

:)